Please help solve this problem. A couple plans to save for their childs' college education. What principal must be deposited when their child is born in order to have $100,000 when the child turns 18? Assume the money earns 4% interst compounded quarterly.
Answer by stanbon(75887)

What principal must be deposited when their child is born in order to have $100,000 when the child turns 18? Assume the money earns 4% interst compounded quarterly.
-------
A(t) = P(1+(r/n))^(nt)
-----
100,000 = P(1+(0.04/4))^(4*18)
----
100,000 = P(1.01)^(72)
----
100,000 = P(2.0471)
P = $48,849.61
=======================
Cheers,
Stan H.
=======================